### v1

# Machine Unlearning in Speech Emotion Recognition via Forget Set Alone

Abstract
Speech emotion recognition aims to identify emotional states from speech signals and has been widely applied in human-computer interaction, education, healthcare, and many other fields. However, since speech data contain rich sensitive information, partial data can be required to be deleted by speakers due to privacy concerns. Current machine unlearning approaches largely depend on data beyond the samples to be forgotten. However, this reliance poses challenges when data redistribution is restricted and demands substantial computational resources in the context of big data. We propose a novel adversarial-attack-based approach that fine-tunes a pre-trained speech emotion recognition model using only the data to be forgotten. The experimental results demonstrate that the proposed approach can effectively remove the knowledge of the data to be forgotten from the model, while preserving high model performance on the test set for emotion recognition.
Index Terms—  Speech emotion recognition, machine unlearning, privacy, adversarial attacks.
1 Introduction
Human speech is an information-rich resource that can provide valuable insights into paralinguistic cues such as emotional states. The growing trend towards more naturalistic human-machine interactions has made the ability to automatically understand and interpret emotions from speech more relevant than ever [1]. Speech Emotion Recognition (SER) has been proposed to automatically identify emotional states from human speech using machine learning methods [2]. SER is promising for manifold applications, such as car-driving [3], education [4], healthcare [5], etc. More recently, end-to-end models, e. g., Wav2Vec [6] and HuBERT [7], trained with self-supervised learning on large-scale datasets have demonstrated a strong capability in extracting abstract emotion-relevant representations. Therefore, they can yield superior performance for SER when fine-tuned on emotional speech datasets. Such good performance promotes the applications of SER based on streaming speech data from various devices, including online platforms, wearable devices, and many others.
Nevertheless, the storage of speech data across multiple platforms and its use in various SER applications can elevate the risk of privacy leakage [8]. Particularly, speech contains a variety of sensitive information usable for identifying the speakers, and inferring their emotions and mental health [9, 10]. Leakage of such sensitive information can cause malicious usages and attacks. For instance, leakage of personal information, e. g., gender and demographic information, can cause the attacks to reduce the model performance in depression detection [11]. In this context, users can request to delete partial speech data to protect their privacy. Meanwhile, even after the data is deleted, SER models still retain information derived from it. Therefore, it is crucial to effectively eliminate the knowledge that these models have learnt from the data.
Machine unlearning has been proposed to train machine learning models for forgetting sensitive data samples, classes, and attributes from a pre-trained model with knowledge of a full dataset [12]. Most machine unlearning approaches are model-agnostic to increase their generability for different SER models. Typical machine unlearning requires both the data to be erased (i. e., forget set) and the remaining data (i. e., remain set) in unlearning. Such a way can maintain the model performance on the original test data for SER. However, leveraging the remain set becomes challenging when other users restrict data redistribution or when the data has already been removed from storage. Additionally, using the remain set is expensive in storage and computing resources in the context of the large volume of speech streams nowadays.
We propose applying a machine unlearning approach using only the data to be forgotten for SER. The proposed approach can (i) train a model to forget the data to be erased, and (ii) maintain model performance using generated adversarial samples that capture the emotional class characteristics of the remain set. The experimental results demonstrate that the proposed approach can train an SER model to forget the erased data, resulting in a reduction of SER accuracy to approx. $0.0\%$ for erased data, while still achieving adequate performance on the original test set.
Related Work. While machine unlearning concepts have been primarily developed for image-related tasks, only a few studies have explored their application in the speech processing domain. Machine unlearning in speech-related tasks mainly focuses on speaker attribute unlearning and instance unlearning. For speaker attribute unlearning, the study in [13] employed domain-adversarial training to identify gender-based violence victim condition and forget speaker identification information. For instance unlearning, the negative gradient method, where the gradient direction is reversed to make the model forget selected samples, was found to be very efficient in a speech recognition task [14]. The study in [15] employed the remaining data during unlearning and proposed a forgetting strategy based on curriculum learning to dynamically learn sample weights [15]. However, as aforementioned, including the remain set can improve the model performance on the original test set, while it may increase the privacy risk of the remain set and require a large storage resource.
For SER, the study in [12] proposed a weight averaging method to combine multiple models, each of which is trained on a data shard. Differently, we focus on instance unlearning of one SER model with the data to be forgotten only, avoiding potential privacy risk of using the remain set. Inspired by the study for image classification [16], we maintain the model performance for SER using generated adversarial samples and elastic model weights.
2 Methodology
Fig. 1: The proposed approach of machine unlearning using randomly labelled and adversarial samples. The model weights are selectively updated during training. Finally, the speech samples are misclassified as the generated random labels.
The training data are represented as $D:\{\bm{x},y\}$, where $\bm{x}$ is the speech samples and $y$ denotes the emotion labels. The speech samples to be erased in machine unlearning are denoted as $D_{e}:\{\bm{x_{e}},y_{e}\}$, and the remaining samples are represented as $D_{r}:\{\bm{x_{r}},y_{r}\}$, i. e., $D=D_{e}\cup D_{r}$, $D_{e}\cap D_{r}=\phi$. Given a pre-trained SER model $f$, the proposed machine unlearning method aims to fine-tune $f$ not only to misclassify $D_{e}$, but also to preserve the knowledge from $D_{r}$. As shown in Fig. 1, the misclassification of $D_{e}$ is achieved by random labelling (see Section 2.1). Adversarial attacks are applied to generate adversarial samples for preserving the model knowledge on $D_{r}$. We further employ elastic model weights during unlearning to preserve more knowledge of $D_{r}$ (see Section 2.2).
2.1 Misclassification of Forget Set
To misclassify $D_{e}$, an effective way is to randomly relabel the data with wrong labels that are different from their original true ones [15]. In this regard, we randomly generate wrong labels for data samples in $D_{e}$. This random relabelling procedure leads to $\hat{D}_{e}:\{\bm{x_{e}},\hat{y}_{e}\}$, $\hat{y}^{i}_{e}\neq y^{i}_{e}$, where $i$ means the sample index. The model is then trained on $\hat{D}_{e}$ for misclassification using the loss function:
$\vskip-2.84526pt\mathcal{L_{\mbox{mis}}}=\mathcal{L_{\mbox{CE}}}(f(\bm{x_{e}}),\hat{y}_{e}),$ (1)
where $\mathcal{L_{\mbox{CE}}}$ denotes the cross-entropy loss.
2.2 Preservation of Knowledge from Remain Set
Random labelling in Section 2.1 can make the model focus on misclassifying the forget set. However, the model can also forget its knowledge learnt from the remain set before unlearning. Using the remain set for unlearning has hidden risks of high computing resources and leakage of other speakers’ privacy. In this regard, the following two methods are applied to preserve the model knowledge on the remain set without using it.
2.2.1 Adversarial Attacks
Adversarial attacks have been shown to have a strong attacking capability to make a model misclassify adversarial data with very poor performance [17]. The adversarial data are usually well-designed and human-indistinguishable from the original real data. In addition to attacking a model, adversarial attacks have also been used for data augmentation, which outperformed typical augmentation methods like random noise [17]. Adversarial data was also demonstrated to contain the feature information of the targeted labels in adversarial attacks [18].
The SER model is trained on generated adversarial data rather than the remain set in this work. Herein, we generate adversarial samples using targeted Projected Gradient Descent (PGD) attacks. Specifically, we randomly assign multiple targeted labels $\{\tilde{y}^{i(1)}_{e},\tilde{y}^{i(2)}_{e},...,\tilde{y}^{i(M)}_{e}\}$ for each sample $x^{i}_{e}$ in $D_{e}$, where $M$ is the number of adversarial samples for each sample to be erased and $\tilde{y}^{i(j)}_{e}\neq y^{i}_{e}$. The PGD attack is a strong attack method with an iteration of Fast Gradient Signed Method (FGSM) in $P$ steps [17]. Given a targeted label $\tilde{y}^{i(j)}_{e}$, $x^{\prime i(j)}_{e}$ is firstly computed by adding $x^{i}_{e}$ and a small random noise with values smaller than $\tau$, where $\tau$ is a constant hyperparameter. The adversarial sample $\tilde{x}^{i(j)}$ is then calculated by PGD based on $x^{\prime i(j)}_{e}$. In such a way, the adversarial samples will be different from each other, especially when $\tilde{y}^{i(j)}_{e}=\tilde{y}^{i(k)}_{e}$, $j\neq k$. In the $t$-th iteration step of PGD, FGSM generates the adversarial sample through the model gradient:
$\vskip-2.84526pt\tilde{x}^{i(j)}_{e(t+1)}=\tilde{x}^{i(j)}_{e(t)}+\sigma*\mbox{sign}(\nabla\mathcal{L}(x^{\prime i(j)}_{e(t)},\tilde{y}^{i(j)}_{e}))$, (2)
where $\nabla$ is stands for the gradient, $\tilde{x}^{i(j)}_{e(1)}=x^{\prime i(j)}_{e}$, and $t=\{1,...,P\}$. Finally, we limit the data difference between adversarial data and real data to be small and invisible with $\tilde{x}^{i(j)}_{e(P)}=\mbox{clip}(\tilde{x}^{i(j)}_{e(P)},x^{i(j)}_{e(P)}-\tau,x^{i(j)}_{e(P)}+\tau)$. Given the adversarial data and targeted labels, the model is trained by
$\vskip-2.84526pt\mathcal{L_{\mbox{adv}}}=\mathcal{L_{\mbox{CE}}}(f(\bm{\tilde{x}_{e}}),\tilde{y}_{e}).$ (3)
2.2.2 Elastic Model Weights
Lifelong learning has been demonstrated its effectiveness in preserving the prior knowledge of a model in transfer learning [19]. To further preserve model knowledge learnt from $D_{r}$, the Elastic Weight Consolidation (EWC) [20] is employed to assign high constraints for important model parameters and low constraints for unimportant parameters. The importance of model parameters is calculated with Fisher matrix [21], which is approaching the second derivative of the loss function. Herein, as the model is expected to forget the data to be erased, we calculate the Fisher matrix via the cross-entropy loss between $f(\bm{x_{e}})$ and $\hat{y}_{e}$. Therefore, high constraints are given to parameters important for misclassifying the forget set. The Fisher matrix $F$ is used in the loss function of EWC:
$\vskip-2.84526pt\mathcal{L}_{\mbox{ewc}}=\sum_{k}F_{k}(\theta_{k}-\theta_{k}^{*})^{2},$ (4)
where $\theta$ denotes the parameters of $f$, and $\theta^{*}$ is the model parameters before machine unlearning.
2.2.3 Model Training
To train a model which can forget the data to be erased and also remember the knowledge of the remain set, the loss function is combined by
$\vskip-2.84526pt\mathcal{L}=\lambda_{1}\mathcal{L}_{\mbox{mis}}+\lambda_{2}\mathcal{L}_{\mbox{adv}}+\lambda_{3}\mathcal{L}_{\mbox{ewc}},$ (5)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are the coefficients for the three loss functions, respectively.
3 Experimental Results
3.1 Database
The Database of Elicited Mood in Speech (DEMoS) [22] is used to validate the proposed approach. The DEMoS corpus is an Italian speech dataset recorded from $68$ speakers (23 females and 45 males) with $9,697$ speech samples in total. The $332$ neutral samples are not used in this study for class balance. The other $9,365$ samples are annotated in seven classes, including Anger, Disgust, Fear, Guilt, Happiness, Sadness, and Surprise. The data is split into training ($3,024$ samples), validation ($3,024$ samples), and test sets ($3,317$ samples) in a speaker-independent setting. The detail of the data distribution can be found in [17, 19].
3.2 Experimental Setup
The pre-trained Wav2Vec 2.0 model [6] on the Librispeech corpus [23] is fine-tuned on the speech samples resampled in $16$ kHz for $20$ epochs with an “Adam” optimiser. The learning rate is experimentally set as 3E53E-5, and the batch size is 16. The fine-tuned Wav2Vec 2.0 is then trained in machine unlearning. The number of adversarial samples per sample to be erased is $M=20$, and the total step number in PGD is experimentally set as $P=50$. The loss coefficients are set as $\lambda_{1}=0.1$, $\lambda_{2}=1$, and $\lambda_{3}=1E3$, in order to balance the three loss functions for stable training. Notably, we implement two training settings: (i) training the model on the training set and testing on the validation set, and (ii) training the model on the combination of the training and validation sets, and testing on the test set. Compared to the process in (i), the combination of the training and validation sets in (ii) can train a stronger model to be tested on the test set. Finally, Unweighted Average Recall (UAR) is employed to evaluate the models’ forgetting capability on the forget set $D_{e}$, and the models’ utility on unseen speakers’ data, i. e., validation and test sets.
Fig. 2: Performance of machine unlearning models on the validation set when forgetting NN number of samples. The proposed approaches are compared: adversarial attacks (Adv.) and adversarial attacks + elastic model weights (Adv. + Ela.).
Table 1: Model performance (UAR) on the forget set ($D_{e}$) and the validation/test set when training models on the train(ing) / train(ing)+val(idation) sets. The NN denotes the number of speech samples to be forgotten. Apart from fine-tuning a Wav2Vec 2.0 model, the baselines of remaining-data-involved unlearning (Remain. Unl.) and random labelling (Ran. Lab.) are compared with the proposed approaches using adversarial attacks (Adv.) and adversarial attacks + elastic model weights (Adv. + Ela.). The best performance of the proposed approach and Ran. Lab. are compared with significant tests (*: $p<0.001$ in a one-tailed z-test).
$N=10$ $N=30$ $N=50$ $N=100$
Train Train+Val Train Train+Val Train Train+Val Train Train+Val
Method $D_{e}$ Val $D_{e}$ Test $D_{e}$ Val $D_{e}$ Test $D_{e}$ Val $D_{e}$ Test $D_{e}$ Val $D_{e}$ Test
Fine-tune 0.907 0.897 0.907 0.897 0.907 0.897 0.907 0.897
Machine Unlearning with forget set and remain set
Remain. Unl. [15] 0.286 0.907 0.143 0.902 0.152 0.904 0.486 0.910 0.206 0.911 0.237 0.899 0.092 0.918 0.151 0.911
Machine Unlearning with forget set only
Ran. Lab. [15] 0.286 0.746 0.000 0.634 0.048 0.440 0.071 0.506 0.191 0.546 0.076 0.519 0.183 0.276 0.054 0.482
Adv. 0.000 0.840 0.000 0.856* 0.071 0.773 0.000 0.747* 0.024 0.738 0.030 0.556 0.011 0.398 0.051 0.384
Adv. Ela. 0.000 0.855* 0.000 0.814 0.024 0.796* 0.000 0.731 0.029 0.739* 0.020 0.727* 0.071 0.429* 0.110 0.331
Baselines. In machine unlearning, the training epoch is $15$ and the optimiser setting is the same as that in fine-tuning. The adversarial-attack-based machine unlearning approach is compared to two baselines of machine unlearning. (i) Remaining-data-involved unlearning. The model is trained not only with random labelling, but also on the remain set $D_{r}$, thereby the loss function is $\mathcal{L}_{\mbox{mis}}+\mathcal{L}_{\mbox{CE}}(f(\bm{x_{r}},y_{r}))$ [15]. (ii) Random labelling. The model is trained with the forget set and randomly generated wrong labels using (1[15].
3.3 Ablation Study
We compare the model performance when forgetting different numbers of speech samples that vary from $10$ to $100$ and when using different clip values $\tau$ varying from $0.1$ to $0.7$ in Fig. 2. The proposed unlearning approaches are compared, including the approach using adversarial attacks, and the one with adversarial attacks and elastic model weights. The model performance of UAR is below $0.4$ when $\tau=0.1$. This might be caused by the adversarial samples that are too close to the real samples in $D_{e}$ and cannot preserve the information of the remain set, when $\tau$ is too small. Correspondingly, the model performance also decreases when $\tau$ is too large as $0.7$, since a large $\tau$ can result in the shift in class distribution or outliers.
When comparing the model performance on $N$, the models perform mainly better when $N$ is smaller. This is reasonable as it is more challenging to forget more data. When $N=30,50,100$, the model performance across different $\tau$ values is not stable as those when $N=10$. The reason might be that the generated adversarial samples sometimes have a shift in class distribution compared to the remaining data’s distribution when $N$ is large. When comparing the two proposed unlearning methods, the performance of models with adversarial attacks and elastic model weights is mostly better than the performance of models using adversarial attacks only. This indicates the effectiveness of the lifelong learning method.
3.4 Results Comparison
In Table 1, we select the best results in the multiple settings of $\tau$ from the model performance on the validation set. The remaining-data-involved unlearning can perform comparably with the fine-tuned model when the number of forgotten samples ($N$) varies from $10$ to $100$. This can be expected as the remain set is involved in unlearning. In comparison, the models trained only on the forget set perform worse than remaining-data-involved unlearning, since the amount of the training data decreases in unlearning and the model can forget the knowledge learnt from the remain set. Both remaining-data-involved unlearning and random labelling methods cause model performance on $D_{e}$ higher than the chance level (i. e., $0.143$ for seven-class classification), which means the models cannot completely forget the knowledge learnt from $D_{e}$.
Compared to the baselines, the proposed two approaches using adversarial attacks can make the model forget the knowledge learnt from $D_{e}$. Both approaches perform on $D_{e}$ with UARs not higher than $0.110$. Both approaches also outperform the random labelling method, indicating the effectiveness of adversarial attacks in augmenting the data and simulating the data distribution of the remain set. When comparing the two proposed approaches, using elastic model weights can further improve the model performance. The reason can be that the EWC method regulates the models to only update unimportant model parameters. Finally, the models in the proposed two approaches perform worse when $N$ increases, which is reasonable. We can still see significant improvement of the model performance compared to random labelling when $N=10,30,50$ ($p<0.001$ in a one-tailed z-test).
4 Conclusion and Future Work
This work proposed a machine unlearning approach using adversarial attacks to protect data privacy hidden in emotional speech. The proposed approach utilises the forget set only, and generates adversarial samples to help the model preserve the knowledge learnt from the remain set. The weights of the speech emotion recognition model are updated by considering parameter importance during unlearning, thereby preserving more knowledge of the remain set. The experimental results indicate that the proposed approach can effectively train the model to forget the date to be erased and still perform well on unseen speakers’ data for emotion recognition. In future efforts, we will validate the effectiveness of the approach in various speech emotion recognition models. We will also investigate the approach for forgetting specific emotional classes and sensitive information, such as gender.